jigglypuff314 (jigglypuff314):

what is the derivative of: (4x)^3 * (2x)^6 ?

OpenStudy (anonymous):

y = [(4x)^3]*[(2x)^6] The Product Rule: d(f•g)/dx = f´g + fg´ y´ = [(4x)^3]*(6)(2x)^5 (2) + [(2x)^6]*(3)(4x)^2 (4) y´ = 12[(4x)^3]*[(2x)^5] + 12[(4x)^2]*[(2x)^6] Factor out a 4x from one term and a 2x from the other in order to add them together... y´ = 12(4x)[(4x)^2]*[(2x)^5] + 12(2x)[(4x)^2]*[(2x)^5] y´ = [48x + 24x]*[(4x)^3]*[(2x)^5] y´ = (72x)*[(4x)^3]*[(2x)^5]

jigglypuff314 (jigglypuff314):

thank you1 :D

OpenStudy (anonymous):

No Problem =D